# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v2

@@ 1 Introduction
Text-to-image (T2I) diffusion models Ho et al. (2020); Song et al. (2020a; b); Nichol & Dhariwal (2021); Rombach et al. (2022); Ho & Salimans (2022) have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions Dhariwal & Nichol (2021); Ramesh et al. (2021); Gal et al. (2022); Betker et al. (2023); Ruiz et al. (2023); Podell et al. (2023); Esser et al. (2024). Alongside these advancements, concerns have also been raised regarding copyright violations Cui et al. (2023); Shan et al. (2023), offensive content Schramowski et al. (2023); Yang et al. (2024b); Zhang et al. (2025), and privacy concerns Carlini et al. (2023); Yang et al. (2023). To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure Kumari et al. (2023); Gandikota et al. (2023); Zhang et al. (2024a). However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality Gandikota et al. (2023); Orgad et al. (2023); Schramowski et al. (2023); Zhang et al. (2024a), compromising the model’s usability. Therefore, beyond ensuring the effective removal of target concepts (i.e., erasure efficacy), concept erasure should also maintain the original semantics of non-target concepts (i.e., prior preservation).
@@
In this light, we introduce Prior Knowledge Refinement, a suite of techniques that strategically and selectively refine the retain set to mitigate the semantic degradation in searching for the null space. Particularly, we propose Influence-based Prior Filtering (IPF), which first quantifies the influence of concept erasure on each non-target concept. It then prunes the retain set by removing minimally affected concepts, preventing the correlation matrix from approaching full rank and thus maintaining an accurate null space. Subsequently, to further enhance prior preservation over the resulting retain set, we propose Directed Prior Augmentation (DPA), which expands the retain set with directed, semantically consistent perturbations to improve retain coverage. In addition, we incorporate Invariant Equality Constraints (IEC) to preserve specific representations, such as the [SOT] token, that should remain unchanged during editing. IEC enforces equality constraints on such invariants to regularize the retaining of essential generation properties. We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks. Overall, our contributions can be summarized as follows:
@@ 2 Related Works
Concept erasure. Current T2I diffusion models inevitably involve unauthorized and offensive generations due to the noisy training data from web Schuhmann et al. (2021; 2022). Apart from applying additional filters or safety checkers Rando et al. (2022); Betker et al. (2023); Rao (2023), prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms. The training-based paradigm fine-tunes model parameters with specific erasure objectives Kumari et al. (2023); Gandikota et al. (2023); Zhang et al. (2024a); Zhao et al. (2024b); Huang et al. (2024); Kim et al. (2024); Zhang et al. (2024b); Zhao et al. (2024a) and additional regularization terms Kumari et al. (2023); Lyu et al. (2024); Lu et al. (2024a). In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure Orgad et al. (2023); Gandikota et al. (2024); Gong et al. (2025). These methods can erase numerous concepts within seconds, demonstrating superior efficiency in practice. Beyond parameter modification, non-parametric methods (e.g., external modules and sampling interventions) have also been explored Schramowski et al. (2023); Wang et al. (2024b); Yoon et al. (2024); Jain et al. (2024); Lee et al. (2025b; a), but they are fragile in open-source settings.
@@ 4.1 Influence-Based Prior Filtering (IPF)
$\bm{\Delta}_{\text{erase}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\|^{2}}_{\text{regularization}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (5)
@@ 4.3 Invariant Equality Constraints (IEC)
In parallel, we identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with varying sampling prompts. One such invariant is the CLIP-encoded [SOT] token. Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, its embedding consistently remains unchanged during T2I process. Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance Ho & Salimans (2022), which also remains unchanged despite prompt variations. Given the invariance of these embeddings, we consider additional protection measures to ensure their outputs remain unchanged during concept erasure. Specifically, we introduce explicit equality constraints over invariants based on Eq. 3:
@@ 5 Experiments
In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure (Appx. D.4), validating our superior prior preservation. The compared baselines include ConAbl Kumari et al. (2023), MACE Lu et al. (2024a), RECE Gong et al. (2025), and UCE Gandikota et al. (2024), which have achieved SOTA performance across various concept erasure tasks. In implementation, we conduct all experiments on SDv1.4 AI (2022) and generate each image using DPM-solver sampler Lu et al. (2022) over 20 sampling steps with classifier-free guidance Ho & Salimans (2022) of 7.5. More implementation details and compared baselines can be found in Appx. C and Appx. D.3.
@@ 5.1 On Few-Concept Erasure
Table 1: Quantitative comparison of the few-concept erasure on instances (left) and artistic styles (right) following Lyu et al. (2024). Arrows indicate the preferred direction for each metric, and the best results are highlighted in bold. Our method consistently improves prior preservation for non-target and general concepts from MS-COCO (shaded in pink) while achieving effective concept erasure. While our CS is not the lowest for target concpet, Appx. D.1 and Fig. 7 show our method is sufficient for erasure, and lower CS may further compromise prior preservation.
@@
Analysis and discussion. Table 1 compares the results of erasing various instance concepts and artistic styles. Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content. Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig. 4 and Appx. D.1. In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge. Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑). A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. UCE’s 10.29 ↑), showing our robustness in erasing increasing target concepts.
@@ 5.2 On Multi-Concept Erasure
Table 2: Quantitative comparison of the multi-concept erasure in erasing 10, 50, and 100 celebrities. The best results are highlighted in bold. Our method can effectively erase up to 100 celebrities simultaneously, achieving low $\text{Acc}_{e}$ (%) and high $\text{Acc}_{r}$ (%) that preserve non-target celebrities with minimal appearance changes. This yields the best overall erasure performance $H_{o}$ and competitive runtime (s) on one A100 GPU, successfully erasing 100 concepts in just 5 seconds.
@@ 5.3 Further Analysis
More applications on other T2I models. To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper Lykon (2023) and RealisticVision SG161222 (2023) from Fig 6 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column). (2) Knowledge editing on SDXL Podell et al. (2023) from Fig 6 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge. Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images. (3) Instance erasure on SDv3 Esser et al. (2024) from Fig 6 (c): To accommodate the diffusion transformer (DiT) Peebles & Xie (2023) architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
@@ 6 Conclusion
We have made extensive efforts to ensure reproducibility of our work. The proposed method, SPEED, is thoroughly described in the main paper (Secs. 3 and 4), with complete theoretical derivations provided in Appx. B. Implementation details, including experimental setup details and erasure configurations, are given in Appx. C. The experimental setups for all three erasure tasks (few-concept, multi-concept, and implicit concept erasure) are described in detail, with complete quantitative results and ablation studies reported in Sec. 5 and Appx. D.
@@ Appendix A Preliminaries
T2I diffusion models. T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) Rombach et al. (2022). Unlike pixel-space diffusion, LDMs operate in the latent space of a pretrained autoencoder, reducing computational costs while maintaining high-quality synthesis. LDMs consist of a vector-quantized autoencoder Van Den Oord et al. (2017); Esser et al. (2021) and a diffusion model Dhariwal & Nichol (2021); Ho et al. (2020); Sohl-Dickstein et al. (2015); Kingma et al. (2021); Song et al. (2020b). The autoencoder encodes an image $\bm{x}$ into a latent representation $\bm{z}=\mathcal{E}(\bm{x})$ and reconstructs it via $\bm{x}\approx\mathcal{D}(\bm{z})$. The diffusion model learns to generate latent codes through a denoising process. The training objective is given by Ho et al. (2020); Rombach et al. (2022):
@@ C.1 Experimental Setup Details
Multi-concept erasure. We then compare methods on multi-concept erasure, a more challenging and realistic concept erasure task. Following the experiment setup from Lu et al. (2024a), we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 AI (2022) can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) Hasty et al. (2019). This dataset is divided into two groups: an erasure set with 10, 50, and 100 celebrities and a retain set with 100 other celebrities. The full list for both sets is presented in Table 4. We experiment with erasing 10, 50, and 100 celebrities with the predefined target concepts and the entire retain set is utilized in all cases. In evaluation, we prepare five celebrity templates, (i.e., “a portrait of {Celebrity}”, “a sketch of {Celebrity}”, “an oil painting of {Celebrity}”, “{Celebrity} in an official photo”, and “an image capturing {Celebrity} at a public event”) and generate 500 images for both sets. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates).
@@ C.2 Erasure Configurations
Implementation of SPEED. In line with previous methods Kumari et al. (2023); Lu et al. (2024a); Gong et al. (2025); Gandikota et al. (2024), we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment Hertz et al. (2022). In contrast, we only edit the value matrices in the CA layers, as suggested by Wang et al. (2024b). This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images Tewel et al. (2023). In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig. 4 and 5 (further ablation comparison is provided in Appx. D.5). The augmentation times $N_{A}$ in Eq. 9 is set to 10 and the augmentation ranks $r$ in Eq. 7 is set to 1 as ablated in Appx. D.5. Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the singular vectors corresponding to the singular values below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following Fang et al. (2024). Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq. 12 to ensure invertibility.

### v4

@@ 1 Introduction
Text-to-image (T2I) diffusion models Ho et al. (2020); Song et al. (2020a; b); Nichol and Dhariwal (2021); Rombach et al. (2022); Ho and Salimans (2022) have facilitated significant breakthroughs in generating highly realistic and contextually consistent images simply from textual descriptions Dhariwal and Nichol (2021); Ramesh et al. (2021); Gal et al. (2022); Betker et al. (2023); Ruiz et al. (2023); Podell et al. (2023); Esser et al. (2024). Alongside these advancements, concerns have also been raised regarding copyright violations Cui et al. (2023); Shan et al. (2023); Yan et al. (2024), offensive content Schramowski et al. (2023); Yang et al. (2024b); Zhang et al. (2025), and privacy concerns Carlini et al. (2023); Yang et al. (2023). To mitigate ethical and legal risks in generation, it is often necessary to prevent the model from generating certain concepts, a process termed concept erasure Kumari et al. (2023); Gandikota et al. (2023); Zhang et al. (2024a). However, removing target concepts without carefully preserving the semantics of non-target concepts can introduce unintended artifacts, distortions, and degraded image quality Gandikota et al. (2023); Orgad et al. (2023); Schramowski et al. (2023); Zhang et al. (2024a), compromising the model’s usability. Therefore, beyond ensuring the effective removal of target concepts (i.e., erasure efficacy), concept erasure should also maintain the original semantics of non-target concepts (i.e., prior preservation).
@@
In this light, we introduce Prior Knowledge Refinement, a suite of techniques designed to strategically and selectively refine the retain set so as to mitigate the semantic degradation in searching for the null space. Particularly, we propose Influence-based Prior Filtering (IPF), which first quantifies the influence of concept erasure on each non-target concept. It then prunes the retain set by removing minimally affected concepts, preventing the correlation matrix from approaching full rank and thus maintaining an accurate null space. Subsequently, to further enhance prior preservation over the resulting retain set, we propose Directed Prior Augmentation (DPA), which expands the retain set with directed, semantically consistent perturbations to improve retain coverage. In addition, we incorporate Invariant Equality Constraints (IEC), which enforce equality constraints on invariant representations (e.g., [SOT] token), to ensure they remain unchanged during editing. We evaluate SPEED on three representative concept erasure tasks, i.e., few-concept, multi-concept, and implicit concept erasure, where it consistently exhibits superior prior preservation across all erasure tasks. Our contributions can be summarized as follows:
@@ 2 Related Works
Concept erasure. Current T2I diffusion models inevitably involve unauthorized and offensive generations due to the noisy training data from web Schuhmann et al. (2021; 2022). Apart from applying additional filters or safety checkers Rando et al. (2022); Betker et al. (2023); Rao (2023), prevailing methods modify diffusion model parameters to erase specific target concepts, mainly categorized into two paradigms. The training-based paradigm fine-tunes model parameters with specific erasure objectives Kumari et al. (2023); Gandikota et al. (2023); Zhang et al. (2024a); Zhao et al. (2024b); Huang et al. (2024); Kim et al. (2024); Zhang et al. (2024b); Zhao et al. (2024a); Wu et al. (2025a) and additional regularization In contrast, the editing-based paradigm edits model parameters using a closed-form solution to facilitate efficiency in concept erasure Orgad et al. (2023); Gandikota et al. (2024); Gong et al. (2025). These methods can erase numerous concepts within seconds, demonstrating superior efficiency in practice. Beyond parameter modification, non-parametric methods (e.g., external modules and sampling interventions) have also been explored Schramowski et al. (2023); Wang et al. (2024b); Yoon et al. (2024); Jain et al. (2024); Lee et al. (2025b; a), but they are fragile in open-source settings.
@@ 4.1 Influence-Based Prior Filtering (IPF)
$\bm{\Delta}_{\text{erase}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\|^{2}}_{\text{regularization}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1},$ (5)
@@ 4.3 Invariant Equality Constraints (IEC)
In parallel, we identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with varying sampling prompts. One such invariant is the CLIP-encoded [SOT] token. Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, its embedding consistently remains unchanged during T2I process. Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance Ho and Salimans (2022), which also remains unchanged despite prompt variations. Given the invariance of these embeddings, we consider additional protection measures to ensure their outputs remain unchanged during concept erasure. Specifically, we introduce explicit equality constraints over invariants based on Eq. 3:
@@ 5 Experiments
In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure (Appx. D.4), validating our superior prior preservation. The compared baselines include ConAbl Kumari et al. (2023), MACE Lu et al. (2024a), RECE Gong et al. (2025), and UCE Gandikota et al. (2024), which have achieved SOTA performance across various concept erasure tasks. In implementation, we conduct all experiments on SDv1.4 Stability AI (2022) and generate each image using DPM-solver sampler Lu et al. (2022) over 20 sampling steps with classifier-free guidance Ho and Salimans (2022) of 7.5. More implementation details and compared baselines can be found in Appx. C and Appx. D.3.
@@ 5.1 On Few-Concept Erasure
Table 1: Quantitative comparison of the few-concept erasure on instances (left) and artistic styles (right), where target concepts are highlighted in pink. Arrows indicate the preferred direction for each metric, and the best results are highlighted in bold. Our method consistently improves prior preservation on non-target and general concepts from MS-COCO while achieving effective concept erasure. Although our CS is not always the lowest for target concepts, Appx. D.1 shows our method is sufficient for erasure, and lower CS may further compromise prior preservation.
@@
Analysis and discussion. Table 1 compares the results of erasing various instance concepts and artistic styles. Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content. Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Figs. 4 and 7. On this basis, lower CS values typically indicate “over-erasure” of the target concept, since further reductions in CS after successful erasure usually come at the cost of prior preservation, as detailed in Appx. D.1. Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑). A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. UCE’s 10.29 ↑), showing our superiority of prior preservation in erasing increasing target concepts.
@@ 5.2 On Multi-Concept Erasure
Table 2: Quantitative comparison of the multi-concept erasure in erasing 10, 50, and 100 celebrities. The best results are highlighted in bold. Our method can effectively erase up to 100 celebrities simultaneously, achieving low $\text{Acc}_{e}$ (%) and high $\text{Acc}_{r}$ (%) that preserve non-target celebrities with minimal appearance changes. This yields the best overall erasure performance $H_{o}$ and competitive runtime (s), successfully erasing 100 concepts in just 5 seconds.
@@ 5.3 Further Analysis
More applications on other T2I models. To validate the transferability of our method across versatile applications, we conduct further experiments on various T2I models with different weights and architectures, including: (1) Composite concept erasure on DreamShaper Lykon (2023) and RealisticVision SG161222 (2023) from Fig 6 (a): Our method can precisely erase the target concept(s) while preserving other non-target elements within the prompt, such as the Van Gogh-style background (2nd column) and the Snoopy character (3rd column). (2) Knowledge editing on SDXL Podell et al. (2023) from Fig 6 (b): The arbitrary nature of anchor concepts allows us to edit the pre-trained model knowledge. Herein, our method effectively edits the model knowledge while maintaining the overall layout and semantics of the generated images. (3) Instance erasure on SDv3 Esser et al. (2024) from Fig 6 (c): To accommodate the diffusion transformer (DiT) Peebles and Xie (2023) architecture in T2I models, we adapt our method to a DiT-based model, demonstrating a well-balanced trade-off between erasure (1st row) and preservation (2nd row) as well.
@@ 6 Conclusion
We have made extensive efforts to ensure the reproducibility of our work. The proposed method, SPEED, is thoroughly described in the main paper (Secs. 3 and 4), with complete theoretical derivations provided in Appx. B. Implementation details, including experimental setup details and erasure configurations, are given in Appx. C. The experimental setups for all three erasure tasks (few-concept, multi-concept, and implicit concept erasure) are described in detail, with complete quantitative results and ablation studies reported in Sec. 5 and Appx. D.
@@ Appendix A Preliminaries
T2I diffusion models. T2I generation has seen significant advancements with diffusion models, particularly Latent Diffusion Models (LDMs) Rombach et al. (2022). Unlike pixel-space diffusion, LDMs operate in the latent space of a pretrained autoencoder, reducing computational costs while maintaining high-quality synthesis. LDMs consist of a vector-quantized autoencoder Van Den Oord et al. (2017); Esser et al. (2021) and a diffusion model Dhariwal and Nichol (2021); Ho et al. (2020); Sohl-Dickstein et al. (2015); Kingma et al. (2021); Song et al. (2020b); Wang et al. (2024c); Li et al. (2025). The autoencoder encodes an image $\bm{x}$ into a latent representation $\bm{z}=\mathcal{E}(\bm{x})$ and reconstructs it via $\bm{x}\approx\mathcal{D}(\bm{z})$. The diffusion model learns to generate latent codes through a denoising process. The training objective is given by Ho et al. (2020); Rombach et al. (2022):
@@ C.1 Experimental Setup Details
Multi-concept erasure. We then compare methods on multi-concept erasure, a more challenging and realistic concept erasure task. Following the experiment setup from Lu et al. (2024a), we introduce a dataset consisting of 200 celebrities, where their portraits generated by SDv1.4 Stability AI (2022) can be recognizable with exceptional accuracy by the GIPHY Celebrity Detector (GCD) Hasty et al. (2019). This dataset is divided into two groups: an erasure set with 10, 50, and 100 celebrities and a retain set with 100 other celebrities. The full list for both sets is presented in Table 4. We experiment with erasing 10, 50, and 100 celebrities with the predefined target concepts and the entire retain set is utilized in all cases. In evaluation, we prepare five celebrity templates, (i.e., “a portrait of {Celebrity}”, “a sketch of {Celebrity}”, “an oil painting of {Celebrity}”, “{Celebrity} in an official photo”, and “an image capturing {Celebrity} at a public event”) and generate 500 images for both sets. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates).
@@ C.2 Erasure Configurations
Implementation of SPEED. In line with previous methods Kumari et al. (2023); Lu et al. (2024a); Wu et al. (2025b); Gong et al. (2025); Gandikota et al. (2024), we edit the cross-attention (CA) layers within the diffusion model due to their role in text-image alignment Hertz et al. (2022). In contrast, we only edit the value matrices in the CA layers, as suggested by Wang et al. (2024b). This choice is grounded in the observation that the keys in CA layers typically govern the layout and compositional structure of the attention map, while the values control the content and visual appearance of the images Tewel et al. (2023). In the context of concept erasure, our goal is to effectively remove the semantics of the target concept, and we find that only editing the value matrices is sufficient as shown in Fig. 4 and 5 (further ablation comparison is provided in Appx. D.5). The augmentation times $N_{A}$ in Eq. 9 is set to 10 and the augmentation ranks $r$ in Eq. 7 is set to 1 as ablated in Appx. D.5. Meanwhile, given that eigenvalues are rarely strictly zero in practical applications when determining the null space, we select the singular vectors corresponding to the singular values below $10^{-1}$ on few-concept and implicit concept erasure and $10^{-4}$ on multi-concept erasure following Fang et al. (2024). Moreover, since the retain set only includes ‘ ’ in implicit concept erasure, we add an identity matrix $\mathbf{I}$ with weight $\lambda=0.5$ to the term $(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2})^{-1}$ in Eq. 12 to ensure invertibility.
Figure 8: Comparison of CS values across different erasure methods. We compare the results in erasing Snoopy and Mickey, and highlight the corresponding CS in blue. Our method achieves successful concept erasure with moderate CS values. In contrast, RECE achieves the lowest CS by enabling more aggressive erasure. For example, removing Snoopy into a landscape without a subject, and changing Mickey into a generic person. We argue that such over-erasure unnecessarily compromises prior preservation as evidenced by Tables 1 and 2.
Appendix D Additional Experiments
D.1 More Demonstrations
Table 5: Human study of erasure efficacy in erasing Snoopy, Mickey, and Spongebob from Table 1.
Snoopy Mickey Spongebob Average
RECE 98.93% 98.27% 99.07% 98.76%
Ours 98.20% 98.40% 98.80% 98.47%
We also evaluate the CS value before and after concept erasure to assess the erasure efficacy. As shown in Fig. 8, our method achieves successful erasure of specific concepts such as Snoopy and Mickey while maintaining moderate CS values (24.18 and 23.44, respectively). This indicates that effective erasure does not require minimizing CS to an extreme. In contrast, RECE obtains the lowest CS (19.79 and 18.75), but this is achieved at the cost of overly aggressive erasure. For example, transforming Snoopy into an unrecognizable image and replacing Mickey with a generic human figure. While such strategies may enhance erasure efficacy, they also risk compromising prior knowledge. This trade-off is reflected in higher non-target FIDs, as shown in Tables 1 and 2.
To further demonstrate that our current erasure is adequate, we additionally conduct a human study. For our method and RECE, we randomly sample 50 generated images per method to erase Snoopy, Mickey, and Spongebob. We then recruit 30 human participants through Amazon Mechanical Turk Amazon to vote “yes” or “no” on whether the target concept is visually erased or not. As shown in Table 5, the overall results (RECE’s 98.76% v.s. Our 98.47%) indicate that our method achieves successful erasure on par with RECE from the human perspective.
D.2 Complete Results on Few-Concept Erasure